# Towards AI as Colleagues: Multi-Agent System Improves Structured Professional Ideation

### v1

@@ Abstract.
Most AI systems today are designed to manage tasks and execute predefined steps. This makes them effective for process coordination but limited in their ability to engage in joint problem-solving with humans or contribute new ideas. We introduce MultiColleagues, a multi-agent conversational system that shows how AI agents can act as colleagues by conversing with each other, sharing new ideas, and actively involving users in collaborative ideation. In a within-subjects study with 20 participants, we compared MultiColleagues to a single-agent baseline. Results show that MultiColleagues fostered stronger perceptions of social presence, produced ideas rated significantly higher in quality and novelty, and encouraged deeper elaboration. These findings demonstrate the potential of AI agents to move beyond process partners toward colleagues that share intent, strengthen group dynamics, and collaborate with humans to advance ideas.
Figure 1. MultiColleagues: An Illustrative User Scenario of AI-Supported Ideation. Example of P10 engages in an ideation session with three role-differentiated AI colleagues. In Step 1, the challenge is introduced. In Step 2, P10 selects colleagues (User Researcher, Data Scientist, and Software Engineer). In Step 3, the discussion unfolds: in Explore mode, the team generates diverse ideas; in Focus mode, they converge on Trustworthy mood-aware Karaoke that balances user experience with privacy, simplicity, and local data processing.:
@@ 1. Introduction
Recent adoption of large language models (LLMs) has moved from deployments as “copilot” tools toward more dynamic roles in collaborative settings. Early applications positioned LLMs as tools or judges (Brown et al., 2020; Ouyang et al., 2022), automating tasks such as summarization, programming assistance, or quality assessment that humans found repetitive or peripheral (Mastropaolo et al., 2023; Mozannar et al., 2024). Although useful, these applications positioned artificial intelligence (AI) as automation rather than as a partner in collaboration (Lim et al., 2009), whereas emerging multi-agent frameworks suggest a broader transition (Dubey et al., 2020; Bansal et al., 2021a). Systems such as AutoGen (Wu et al., 2023) and CrewAI (cre, 2024) demonstrate how multiple LLMs can be coordinated under structured protocols, with agents adopting complementary roles such as planner, critic, or explainer. LLMs are moving beyond single-task execution to acting as participants that contribute perspectives in a team process.
@@
The growing use of multi-agent frameworks prompts a deeper question of whether LLMs can be experienced not only as tools but as colleagues in collaborative work. Collaboration between humans and models is often most productive when it leverages asymmetries in capability, with people contributing judgment, values, and imagination while LLMs provide scale, recall, and breadth (Dellermann et al., 2019; Amershi et al., 2019). Moving beyond copilot metaphors requires examining whether users can experience these systems as team members with complementary strengths.
@@
Ideation provides a representative context to probe this question. Research in creativity shows that human strengths remain decisive in the early stages of idea generation. At the same time, studies show that LLM assistance can expand the number and diversity of ideas, while also risking homogenization and over-reliance (Buçinca et al., 2021; Lu et al., 2024; Padmakumar and He, 2024). Multi-agent personas present a promising approach, as they can emulate interdisciplinary team dynamics, introduce diverse perspectives, and enable humans to retain strategic oversight. Interaction paradigms play a central role in shaping this experience. Roundtable exchanges, hierarchical supervision, or progressive disclosure influence how people engage with multiple agents and how cognitive load is managed (Zamfirescu-Pereira et al., 2023; Luger and Sellen, 2016). Besides, role-playing structures task allocation and heightens social presence, which makes collaboration feel closer to working with teammates (Lee et al., 2019; Ashktorab et al., 2019).
@@
Recent work has examined distinct facets of this design space, with systems emphasizing either divergent–convergent structuring through staged debate and role play (Lu et al., 2024) or orchestration through contrasting roles and rotating perspectives to guide reflection (Zhang et al., 2024; Pang et al., 2025). These approaches demonstrate the value of both dimensions but have largely been explored in isolation. Building on these directions, we introduce MultiColleagues, a multi-agent conversational system that brings together diverse AI personas for co-ideation while centering the human as facilitator-in-chief. The system pursues three design goals: (1) supporting shifts between divergent and convergent thinking, (2) engaging diverse viewpoints to expand the idea space, and (3) providing clear features that preserve human oversight in collaboration.
Guided by these goals, we evaluate MultiColleagues through a within-subjects study against a single-agent ChatGPT baseline, focusing on three research questions:
@@
    RQ1 - Experience: How do role-taking patterns and perceptions of social presence shape the collaborative atmosphere and user engagement?
@@
    RQ2 - Outcomes: How does exposure to multiple AI-Colleague perspectives affect the creative outcomes, and what underlying dynamics may explain these differences?
@@
    RQ3 - System Design: How do system design features support or constrain support during creative ideation?
@@ 2.1. From Tools to Colleagues: The Evolution of Human–AI Collaboration
Over recent years, large language models (LLMs) have emerged as transformative tools across a wide range of applications, demonstrating state-of-the-art performance in natural language processing and knowledge-intensive tasks (Yang et al., 2024; Lappin, 2024). They have been widely adopted in domains such as software development (Wu et al., 2024; Cui, 2024), writing and creativity support (Clark et al., 2022; Wu et al., 2025), scientific research and literature review (Menick et al., 2022; Gao and other collaborators, 2023), and scientific experimentation (Bran et al., 2023). Across these domains, LLMs have primarily functioned as assistants or evaluators, supporting human work by improving efficiency and facilitating decision-making (Qin et al., 2023; Yuan et al., 2022). Nevertheless, human-in-the-loop (HITL) involvement remains indispensable. Research has shown that relying on humans solely as “reviewers” can introduce risks, such as decision-making risks (Schoeffer et al., 2024; Ghosh and Rintel, 2025), reliability and transparency risks (Amirizaniani et al., 2024), systemic risks (Reid et al., 2025), and ethical risks (Peter et al., 2025; Akbulut et al., 2024; Kocaballi et al., 2022). Additional HITL studies across domains reinforce the irreplaceable role of humans and confirm that human involvement enhances accuracy and reliability (Li et al., 2025; Takerngsaksiri et al., 2025). Building on this foundation, human-centered and mixed-initiative frameworks argue that the future of AI lies in augmentation rather than replacement, coupling higher levels of automation with sustained human control (Shneiderman, 2022; Amershi et al., 2019). As LLMs grow more capable, systems are gradually shifting toward proactive collaboration. Multi-agent frameworks such as AutoGen and CAMEL operationalize LLMs as differentiated collaborators (Li et al., 2023a; Wu et al., 2023), where models can assume specific roles, engage in mutual critique, and coordinate planning activities. Through such structured interactions, they approximate team-like collaboration and move toward systems where LLMs function less as tools and more as partners (Du et al., 2023; Liu et al., 2023).
2.2. Multi-Agent LLMs and Teamwork Dynamics
The growing capabilities of recent years’ LLMs have motivated increasing interest in multi-agent frameworks as a way to extend the scope and complexity of applications. A first major direction examines task-decomposed collaboration, where models are assigned complementary roles such as planning (Wu et al., 2023), execution (Wang et al., 2024a), or debugging (Dibia et al., 2024). Within this strand, researchers have proposed different coordination paradigms, including dialogic debate (Du et al., 2023; Liang et al., 2023) and hierarchical supervision (Guo et al., 2024; Wu et al., 2024). While these systems demonstrate gains in reasoning, factuality, and coding ability, their evaluations remain largely confined to benchmark settings (Hendrycks et al., 2020; Cobbe et al., 2021) rather than interactive use.
A second line of work explores persona-driven role play, showing that LLMs can convincingly simulate interdisciplinary teamwork by adopting distinct identities. This stream highlights how personas shape more human-like interaction styles (Park et al., 2023; Li et al., 2023b; Wei et al., 2023; Shao et al., 2023), enhance engagement (Wang et al., 2024b; Cui et al., 2023), and diversify task performance through structured impersonation (Li et al., 2023a). Examples range from predefined expert roles in CAMEL (Li et al., 2023a), which bring complementary perspectives to task execution, to generative agents that exhibit emergent social behaviors in daily scenarios (Park et al., 2023; Shanahan et al., 2023). Collectively, this work points to the potential of role differentiation for strengthening social presence and aligning collaboration with human expectations.
A third line of research focuses on coordination mechanisms that sustain coherence across longer or more complex interactions. Efforts here include shared memory, scheduling, and blended model outputs. Skeleton-guided reasoning (Ning et al., 2023), model fusion (Jiang et al., 2023), and collaborative decoding strategies (Sun et al., 2023) exemplify how multiple agents can combine strengths to tackle tasks beyond the capacity of a single model. At the same time, across all three directions, prior work consistently emphasizes the importance of human oversight for aligning decisions with user intent and preventing failures under high autonomy (Bansal et al., 2021b; Saunders et al., 2017; Amershi et al., 2014; Christiano et al., 2017).
As shown in Table 1, recent work has begun to explore multi-agent systems in interactive settings. LLM Discussion (Lu et al., 2024) adopts a three-phase agent–agent debate with role-play to enhance originality and elaboration of model outputs. SWTW (Zhang et al., 2024) introduces progressively contrasting roles in media reading, using orchestration and gamified puzzles to mitigate filter bubbles and deepen reflection. Weaver (Pang et al., 2025) organizes advisory-style roundtables with speaker rotation and summaries to anticipate broader social impacts. While each contributes valuable orchestration strategies, they do not combine dynamic turn-taking, divergent–convergent phases, and user-facing facilitation within controlled, in-situ studies. Motivated by these gaps, MultiColleagues advances co-ideation by combining dynamic turn-taking, explicit divergent–convergent shifts, and interactive orchestration, enabling a controlled study of how role differentiation and multi-agent dynamics shape collaborative ideation.
@@ 2.2. Multi-Agent LLMs and Teamwork Dynamics
Turn Div–Conv Orch. In-situ Pick Ctrl.
@@
Table 1. Comparison of multi-agent systems. Icons: yes, — not applicable. Turn = dynamic turn selection; Div–Conv = divergent–convergent phases; Orch. = human-facing orchestration; In-situ = interactive study; Pick = user chooses agents; Ctrl. = controlled (within-subjects) vs. baseline.
@@ 2.3. GenAI-Assisted Ideation
Research on AI-assisted ideation has evolved from retrieval-based systems to generative, structured support. Early tools such as IdeaHound (Siangliulue et al., 2016), ProbMap (MacNeil et al., 2021), and IdeateRelate (Xu et al., 2021) used semantic similarity to surface relevant ideas but offered limited scaffolding beyond recall. Recent systems embed generative models into interactive environments: Jamplate organizes reflection through templates (Xu et al., 2024), BioSpark clusters LLM concepts into analogical “inspiration cards” (Kang et al., 2025), Scideator recombines scientific paper facets to propose novel research ideas (Radensky et al., 2025), and CausalMapper visualizes causal relations among concepts to guide systematic exploration (Huang et al., 2023). These systems demonstrate how generative outputs can be transformed into higher-level structures that promote reflection, analogy, and cross-domain thinking. Complementary approaches emphasize role-based scaffolding. PersonaFlow (Liu et al., 2025) simulated multiple expert personas to inject disciplinary perspectives, while Rayan et al. (Rayan et al., 2024) leveraged generative chat to stimulate collaboration. Yet purely textual interfaces risk fixation: listing example solutions can anchor users to narrow trajectories (Chan et al., 2024). Spatial or hierarchical arrangements mitigate this risk by situating ideas in broader solution spaces, as demonstrated by tree-structured diagrams for structured exploration (Ding et al., 2025; Yang et al., 2025).
More recently, attention has shifted to the limitations of single-agent LLMs. Their autoregressive nature tends toward convergence, limiting diversity (Wadinambiarachchi et al., 2024). Multi-agent approaches distribute perspectives across distinct personas to sustain divergence and reframe models as collaborators rather than tools (Fukumura and Ito, 2025; Ghosh and Rintel, 2025). Building on this trajectory, our work positions LLMs as colleagues in a multi-agent environment, embedding the Double Diamond design model (Design Council, 2005) to balance divergence and convergence, user control and agent autonomy, toward more coherent yet diverse ideation.

### v2

# Towards AI as Colleagues: Multi-Agent System Improves Structured Ideation Processes

@@ Abstract.
Most AI systems today are designed to manage tasks and execute predefined steps. This makes them effective for process coordination but limited in their ability to engage in joint problem-solving with humans or contribute new ideas. We introduce MultiColleagues, a multi-agent conversational system that shows how AI agents can act as colleagues by conversing with each other, sharing new ideas, and actively involving users in collaborative ideation processes. In a within-subjects study with 20 participants, we compared MultiColleagues to a single-agent baseline. Results show that MultiColleagues fostered stronger perceived social presence, and participants rated their outcomes as higher in quality and novelty, with more elaboration during ideation. These findings demonstrate the potential of AI agents to move beyond process partners toward colleagues that share intent, strengthen group dynamics, and collaborate with humans to advance ideas.
AI as Colleagues, Multi-Agent Systems, Large Language Models, Interaction Design, Collaborative Ideation
copyright: rightsretainedjournalyear: 2026copyright: ccconference: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems; April 13–17, 2026; Barcelona, Spainbooktitle: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems (CHI ’26), April 13–17, 2026, Barcelona, Spaindoi: 10.1145/3772318.3790375isbn: 979-8-4007-2278-3/2026/04ccs: Human-centered computing Empirical studies in HCIccs: Human-centered computing Collaborative interactionccs: Human-centered computing Empirical studies in collaborative and social computingccs: Human-centered computing Interaction design
Figure 1. MultiColleagues: An Illustrative User Scenario of AI-Supported Ideation. P10 introduces a karaoke-in-autonomous-vehicle challenge (Step 1), selects three AI colleagues (Step 2), then moves from Explore (divergent ideas) to Focus (convergence) on Trustworthy mood-aware Karaoke with privacy, simplicity, and local data processing (Step 3).
Participant 10 scenario with MultiColleagues. The figure depicts Participant 10 identifying a challenge, selecting role-differentiated colleagues, and joining the discussion. Through explore and focus modes, divergent and convergent thinking are applied to generate and refine ideas.
@@ 1. Introduction
Recent adoption of large language models (LLMs) has moved from deployments as “copilot” tools toward more dynamic roles in collaborative settings. Early applications positioned LLMs as tools or judges (Brown et al., 2020; Ouyang et al., 2022), automating tasks such as summarization, programming assistance, or quality assessment that humans found repetitive or peripheral (Mastropaolo et al., 2023; Mozannar et al., 2024). Although useful, these applications positioned artificial intelligence (AI) as automation rather than as a partner in collaboration (Hwang, 2022; Lim et al., 2009), whereas emerging multi-agent frameworks suggest a broader transition (Dubey et al., 2020; Bansal et al., 2021). Systems such as AutoGen (Wu et al., 2023) and CrewAI (25) demonstrate how multiple LLMs can be coordinated under structured protocols, with agents adopting complementary roles such as planner, critic, or explainer. LLMs are moving beyond single-task execution to acting as participants that contribute perspectives in a team process.
@@
The growing use of multi-agent frameworks raises a deeper question: whether LLMs can be experienced not only as tools but also as colleagues in collaborative work. Human–model collaboration is often most productive when it leverages asymmetries in capability, with people contributing judgment, values, and imagination while LLMs provide scale, recall, and breadth (Dellermann et al., 2019; Amershi et al., 2019). Moving beyond copilot metaphors requires examining whether users can experience these systems as team members with complementary strengths.
@@
Ideation provides a representative context to probe this question. Research in creativity shows that human strengths remain decisive in the early stages of idea generation. At the same time, studies show that LLM assistance can expand the number and diversity of ideas, while also risking homogenization and over-reliance (Buçinca et al., 2021; Lu et al., 2024). Multi-agent personas present a promising approach, as they can emulate interdisciplinary team dynamics, introduce diverse perspectives, and enable humans to retain strategic oversight. Interaction paradigms play a central role in shaping this experience. Roundtable exchanges, hierarchical supervision, or progressive disclosure influence how people engage with multiple agents and how cognitive load is managed (Zamfirescu-Pereira et al., 2023; Luger and Sellen, 2016). Besides, role-playing structures task allocation and heightens social presence, which makes collaboration feel closer to working with teammates (Lee et al., 2019; Ashktorab et al., 2019).
@@
Recent work has examined distinct facets of this design space (Table 1), with systems emphasizing divergent and convergent structuring through staged debate and role play (Lu et al., 2024; Chen et al., 2025b), double diamond–inspired workflows (Heyman et al., 2024), and role-based orchestration that rotates perspectives to support reflection (Zhang et al., 2024; Pang et al., 2025). These approaches highlight different core elements of collaborative ideation but have largely been explored in separation. However, real-world brainstorming blends these elements as teams shift between divergent and convergent thinking, draw on multiple viewpoints, and rely on facilitation to stay aligned. This gap between fragmented system mechanisms and integrated human practice motivates systems that better reflect how collaborative ideation unfolds. MultiColleagues addresses this gap by integrating these mechanisms into a unified multi-agent system. Unlike workflow-based systems which structure idea generation through predefined stages (Heyman et al., 2024; Chen et al., 2025b), MultiColleagues models team cognition by having AI personas build on each other’s reasoning, surface differences, and adapt through dynamic turn-taking and facilitator oversight. Building on these approach, we introduce MultiColleagues, a multi-agent conversational system that brings together diverse AI personas for co-ideation while centering the human as facilitator-in-chief. The system follows three design goals: (1) supporting shifts between divergent and convergent thinking, (2) engaging diverse viewpoints to expand the idea space, and (3) providing clear features that preserve human oversight in collaboration.
To examine the value of this social orchestration paradigm, we conducted a within-subjects study that compares MultiColleagues with a state-of-the-art single agent system that reflects the common AI as tool model in ideation practice. This contrast provides a clear basis for assessing whether an AI as colleague design changes how users collaborate, explore ideas, and perceive the interaction. Guided by this comparison, we investigate three research questions:
@@
    RQ 1: How do role-taking patterns and perceptions of social presence shape the collaborative atmosphere and user engagement?
@@
    RQ 2: How does exposure to multiple AI-Colleague perspectives influence users’ ideation processes and perceived outcome quality and novelty?
@@
    RQ 3: How do system design features support or constrain support during creative ideation?
@@ 2. Related Work
Table 1. Comparison of multi-agent systems. Icons: yes, — not applicable. Turn = dynamic turn selection; Div–Conv = divergent–convergent phases; Orch. = human-facing orchestration; Pers. = presence of explicit agent identities or role styles; In-situ = interactive study; Pick = user chooses agents; Ctrl. = controlled (within-subjects) vs. baseline, with indicating comparison against ChatGPT.
@@
Turn Div–Conv Orch. Pers. In-situ Pick Ctrl.
@@
Supermind (CI’24) (Heyman et al., 2024) Single agent tool using double-diamond inspired moves for guided exploration.
CoExploreDS (CHI’25) (Chen et al., 2025b) Design space tool using a div-conv approach for structured exploration.
@@ 2.1. From Tools to Colleagues: The Evolution of Human–AI Collaboration
Over recent years, large language models (LLMs) have emerged as transformative tools across a wide range of applications, demonstrating state-of-the-art performance in natural language processing and knowledge-intensive tasks (Yang et al., 2024; Lappin, 2024). They have been widely adopted in domains such as software development (Wu et al., 2024; Cui et al., 2024), writing and creativity support (Clark et al., 2022; Wu et al., 2025) and scientific experimentation (Bran et al., 2023). Across these domains, LLMs have primarily functioned as assistants or evaluators, supporting human work by improving efficiency and facilitating decision-making (Qin et al., 2023; Yuan et al., 2022). Nevertheless, human-in-the-loop (HITL) involvement remains indispensable. Research has shown that relying on humans solely as “reviewers” can introduce risks, such as decision-making risks (Schoeffer et al., 2024; Ghosh and Rintel, 2025), reliability and transparency risks (Amirizaniani et al., 2024), systemic risks (Reid et al., 2025), and ethical risks (Peter et al., 2025; Akbulut et al., 2024; Kocaballi et al., 2022). Additional HITL studies across domains reinforce the irreplaceable role of humans and confirm that human involvement enhances accuracy and reliability (Li et al., 2025; Takerngsaksiri et al., 2024). Building on this foundation, human-centered and mixed-initiative frameworks argue that the future of AI lies in augmentation rather than replacement, coupling higher levels of automation with sustained human control (Shneiderman, 2022; Amershi et al., 2019). As LLMs grow more capable, systems are gradually shifting toward proactive collaboration. Multi-agent frameworks such as AutoGen and CAMEL operationalize LLMs as differentiated collaborators (Li et al., 2023b; Wu et al., 2023), where models can assume specific roles, engage in mutual critique, and coordinate planning activities. Through such structured interactions, they approximate team-like collaboration and move toward systems where LLMs function less as tools and more as partners (Du et al., 2023; Liu et al., 2023).
2.2. Multi-Agent LLMs and Teamwork Dynamics
The growing capabilities of recent years’ LLMs have motivated increasing interest in multi-agent frameworks as a way to extend the scope and complexity of applications. A first major direction examines task-decomposed collaboration, where models are assigned complementary roles such as planning (Wu et al., 2023), execution (Wang et al., 2024a), or debugging (Dibia et al., 2024). Within this strand, researchers have proposed different coordination paradigms, including dialogic debate (Du et al., 2023; Liang et al., 2023) and hierarchical supervision (Guo et al., 2024; Wu et al., 2024). While these systems demonstrate gains in reasoning, factuality, and coding ability, their evaluations remain largely confined to benchmark settings (Hendrycks et al., 2020) rather than interactive use.
System workflow of persona-guided discussions. The figure shows five steps: Step 1, preparing discussion by selecting colleagues; Step 2, system ranks and selects a persona and generates a reply; Step 3, output is presented, awaiting user action; Step 4, user responds, continues, or calls a facilitator; Step 5, all logs are stored in the database.
A second line of work explores persona-driven role play. Recent studies show that AI generated personas are sufficiently coherent to function as stable identities in collaborative settings (Schuller et al., 2024; Smrke et al., 2025; Salminen et al., 2024; Jung et al., 2025; Pang et al., 2025), enabling LLMs to convincingly simulate interdisciplinary teamwork by adopting distinct identities. This stream highlights how personas shape more human-like interaction styles (Park et al., 2023; Li et al., 2023a; Wei et al., 2023; Shao et al., 2023), enhance engagement (Wang et al., 2024b; Cui et al., 2023), and diversify task performance through structured impersonation (Li et al., 2023b). Examples range from predefined expert roles in CAMEL (Li et al., 2023b), which bring complementary perspectives to task execution, to generative agents that exhibit emergent social behaviors in daily scenarios (Park et al., 2023; Shanahan et al., 2023). Collectively, this work points to the potential of role differentiation for strengthening social presence and aligning collaboration with human expectations.
A third line of research focuses on coordination mechanisms that sustain coherence across longer or more complex interactions. Efforts here include shared memory, scheduling, and blended model outputs. Skeleton-guided reasoning (Ning et al., 2023), model fusion (Jiang et al., 2023), and collaborative decoding strategies (Sun et al., 2023) exemplify how multiple agents can combine strengths to tackle tasks beyond the capacity of a single model. At the same time, across all three directions, prior work consistently emphasizes the importance of human oversight for aligning decisions with user intent and preventing failures under high autonomy (Bansal et al., 2021; Amershi et al., 2014; Christiano et al., 2017).
As shown in Table 1, recent work has begun to explore multi-agent systems in interactive settings. LLM Discussion adopts a three-phase agent–agent debate with role-play to enhance originality and elaboration of model outputs (Lu et al., 2024). SWTW introduces progressively contrasting roles in media reading, using orchestration and gamified puzzles to mitigate filter bubbles and deepen reflection (Zhang et al., 2024). Weaver organizes advisory-style roundtables with speaker rotation and summaries to anticipate broader social impacts (Pang et al., 2025). While each contributes valuable orchestration strategies, they do not combine dynamic turn-taking, divergent–convergent phases, and user-facing facilitation within controlled, in-situ studies. Motivated by these gaps, MultiColleagues advances co-ideation by combining dynamic turn-taking, explicit divergent–convergent shifts, and interactive orchestration, enabling a controlled study of how role differentiation and multi-agent dynamics shape collaborative ideation.
@@ 2.3. GenAI-Assisted Ideation
Recent works on AI-assisted ideation spans a wide range of approaches that enrich creative exploration. Early systems emphasized retrieving relevant concepts to spark new ideas, either through retrieval-based pipelines that surfaced related crowd-generated content (Siangliulue et al., 2016) or through similarity-driven semantic mapping that highlighted adjacent concepts (MacNeil et al., 2021; Xu et al., 2021). With the rise of generative models, newer tools embed LLM outputs into interactive structures that support higher-level reasoning, such as template-guided reflection (Xu et al., 2024), analogical inspiration cards derived from LLM-generated concepts (Kang et al., 2025), recombined scientific facets for idea expansion (Radensky et al., 2025), and causal diagrams that help users explore conceptual relationships (Huang et al., 2023). Other systems introduce varied perspectives by incorporating conversational prompting (Rayan et al., 2024) or using role conditioning to simulate expert viewpoints (Liu et al., 2025). Collectively, these works demonstrate the breadth of mechanisms through which generative AI can support creative thinking.
With recent advances in large language models, conversational systems have become effective day to day ideation tools. ChatGPT supports fast concept expansion, iterative refinement, and low effort interaction, making it a widely used single agent workflow for brainstorming and creative exploration (OpenAI, 2024). Reflecting its central role in practice, recent creativity-support systems often build on GPT or benchmark their structured interfaces against ChatGPT-style prompting (Wang et al., 2025b; Heyman et al., 2024; Shen et al., 2025). Despite this widespread adoption, single-agent LLMs remain constrained by their autoregressive nature, which tends toward convergent trajectories (Wadinambiarachchi et al., 2024), and by the limited viewpoint available from a single generative source. These constraints have motivated interest in introducing additional structure or complementary perspectives into the ideation process. Existing work explores individual mechanisms such as divergence–convergence structuring (Lu et al., 2024; Chen et al., 2025b), Double Diamond workflows (Heyman et al., 2024), or modeled perspective shifts (Zhang et al., 2024; Pang et al., 2025). Multi-agent approaches have also emerged as a way to distribute reasoning across personas to encourage more diverse exploration (Fukumura and Ito, 2025; Ghosh and Rintel, 2025). These mechanisms support different parts of the ideation process but remain largely separate, without enabling the coordinated interaction of perspectives and phases typical of real brainstorming. MultiColleagues integrates these elements within a single system, which offers a unified, team-based alternative to a single-agent ChatGPT workflow.